256x^2 = 36
a. -3/8, 3/8
b.-9/64,9/64
c.-8/3,8/3
d. -64/9,64/9

Answer by stanbon(75887)

256x^2 = 36
x^2 = 36/256
x^2 = 9/64
x = 3/8 or x = -3/8
============================
Cheers,
Stan H.